Find the general solution of the following

2x2y''-xy'+(x-5)y=0



1
Expert's answer
2021-08-16T14:02:43-0400

Find the general solution of the following

2x2y''-xy'+(x-5)y=0

Solution:

The general solution of the equation

"x^2y''+axy'+(bx^n+c)y=0", "n\neq 0"

is:

"y=x^\frac{1-a}{2}[C_1J_\nu (\frac2n\sqrt{b}x^\frac n2)+C_2Y_\nu (\frac2n\sqrt{b}x^\frac n2)]" ,

where "\nu=\frac 1n\sqrt{(1-a)^2-4c}" , "C_1" and "C_2" are arbitrary constants;

"J_\nu (z)" and "Y_\nu(z)" are the Bessel functions of the first and second kind.

References :

Kamke, E., Differentialgleichungen: Losungsmethoden und Losungen, I, Gewohnliche Differentialgleichungen, B. G. Teubner, Leipzig, 1977.

Polyanin, A. D. and Zaitsev, V. F., Handbook of Exact Solutions for Ordinary Differential Equations, 2nd Edition, Chapman & Hall/CRC, Boca Raton, 2003.


In our case:

"2x^2y''-xy'+(x-5)y=0"

divided by 2:

"x^2y''-\frac12xy'+(\frac12x-\frac52)y=0"

"a=-\frac12" , "b=\frac12" , "c=-\frac52" , "n=1", "\nu=\sqrt{\frac94+10}=\frac72" .

Solution:

"y=x^\frac{3}{4}[C_1J_\frac72 (\sqrt{\frac{x}{2}})+C_2Y_\frac72 (\sqrt{\frac{x}{2}})]" ,

where "C_1" and "C_2" are arbitrary constants;

"J_\nu (z)" and "Y_\nu(z)" are the Bessel functions of the first and second kind.

Answer: "y=x^\frac{3}{4}[C_1J_\frac72 (\sqrt{\frac{x}{2}})+C_2Y_\frac72 (\sqrt{\frac{x}{2}})]" .